# UNGER: Generative Recommendation with A Unified Code via Semantic and Collaborative Integration

### v2

@@ 4.2. Performance Comparison (RQ1)
Baselines.We compare UNGER with several representative related baselines, encompassing classical sequential modeling approaches as well as the latest emergent generative techniques.
@@
  • ColaRec (Wang et al., 2024a) integrates user-item interactions and content data within an end-to-end framework, leveraging pretrained collaborative identifiers, an item indexing task, and contrastive loss to align semantic and collaborative spaces for enhanced recommendation performance.
@@
  • EAGER (Wang et al., 2024b) integrates behavioral and semantic information through a two-stream architecture with shared encoding, separate decoding, and strategies for contrastive and semantic-guided learning to enhance collaborative information utilization.
@@
Results. Table 3 reports the overall performance of three datasets. The results for all baselines without the superscript are taken from the publicly accessible results (Zhou et al., 2020; Wang et al., 2024b). For missing statistics, we reimplement the baseline and report our experimental results. From the results, we have the following observations:
  • UNGER almost achieves better results than base models among different datasets. Specifically, on the Beauty dataset, compared to the second best baseline EAGER, we achieve improvements of 14.68% and 7.85% in Recall@20 and NDCG@20, respectively, which represents a significant improvement in the recommendation field. We attribute the improvements to UNGER’s effective exploitation of the complementary strengths of collaborative and semantic knowledge with a unified code, which allows it to achieve superior performance relative to EAGER, despite utilizing approximately half the number of total parameters.
  • UNGER beats the previous generative models on most datasets. Compared to previous generative models, UNGER employs a two-stage multi-task joint optimization framework to integrate collaborative and semantic knowledge with a unified code, which successfully captures the interplay between these two types of knowledge, allowing for a more thorough and efficient exploitation of their complementary strengths. The improvements validate the effectiveness of our method and demonstrate the necessity of a unified code.
  • In most cases, generative models outperform classical models. Unlike traditional sequence modeling approaches, which use non-informative item IDs as identifiers, generative models assign hierarchical codes to items as identifiers. This hierarchical structure inherently encodes prior knowledge, facilitating a more effective alignment with the token-by-token generation framework of the Transformer decoder.
@@
Figure 5. T-SNE visualization of different embeddings distribution after applying K-means clustering (clusters=10).
[…]
@@ 4.3. Semantic Domination Issue (RQ2)
To investigate whether our method addresses the semantic dominance issue, we conduct experiments on the Amazon Beauty dataset. We first extract the semantic embeddings of the items, then use only the item IDs from historical interaction data and apply the classical DIN model to obtain the collaborative embeddings. Following the common approach of fusing knowledge from different modalities, we concatenate the semantic embeddings with the collaborative embeddings. Notably, to mitigate the issue caused by the large dimensional differences between embeddings from the two modalities, we first apply PCA to reduce the dimension of the semantic embeddings to match that of the collaborative embeddings, and then concatenate them to obtain the concatenation embeddings. In addition, we obtain the integrated embeddings of items using our proposed method. We use T-SNE to visualize the distribution of these different embeddings. To more clearly illustrate the differences in their distributions, we apply K-means clustering algorithm to these embeddings, using different colors to represent data points from each cluster, as shown in Figure 5.
@@
As shown in Table 4, the concatenation method exhibits a significant disparity in similarity proportions between semantic and collaborative modalities, and Figure 5 further demonstrates that the embeddings obtained by the concatenation method are highly similar to the distribution of semantic embeddings. These observations confirm the presence of the semantic dominance problem. The traditional concatenation method tends to overly emphasize semantic knowledge while greatly ignoring collaborative knowledge, which poses a substantial obstacle to downstream recommendation effectiveness. In contrast, our proposed method effectively addresses this issue. It not only learns embeddings from both collaborative and semantic modalities in a more balanced manner at the distributional level but also achieves a more proportional similarity from each modality in quantitative evaluations. This prevents the semantic embeddings from dominating the final representation. As illustrated in Table 5, our method significantly outperforms both single-modality approaches and the concatenation-based fusion method, achieving the best results across all evaluation metrics. Interestingly, the concatenation method performs worse than the semantic-only approach in Recall@10 and Recall@20, indicating a negative impact on performance. This further highlights the severity of the semantic dominance problem and underscores the advantages of our proposed approach. We attribute these improvements to the successful integration of collaborative and semantic knowledge, which, unlike the concatenation method, better captures the complementary nature of the two different knowledge, enabling a more balanced and effective fusion of information from both modalities.
@@
Figure 6. Impact of the number of model layers on the Beauty dataset. The number of layers varies from 1 to 32.
[…]
@@ 4.4. Ablation Study (RQ3)
We conducted a thorough evaluation of the contribution of each component in UNGER. Specifically, we incrementally removed the intra-modality knowledge distillation (IKD) task and the cross-modality knowledge alignment (CKA) task to create the ablation architectures. The results in Table 6 reveal the following insights:
  • Removing either the CKA or IKD task degrades the model’s performance, while omitting both results in the worst outcome (i.e., the baseline model obtained by simply concatenating semantic and collaborative embeddings). These findings confirm the effectiveness of the two proposed tasks.
  • The introduction of the CKA task enhances recommendation performance, as it effectively integrates knowledge from both collaborative and semantic modalities. By leveraging their complementary characteristics, the CKA task enables the model to learn more expressive and informative item representations.
  • Incorporating both the CKA and IKD tasks allows the model to achieve state-of-the-art performance. Building on the foundation established by the CKA task, the IKD task further promotes the learning of integrated knowledge by encouraging the model to better exploit the complementary strengths of collaborative and semantic knowledge. The observation also highlights the crucial role of both tasks in enabling the model to acquire more powerful unified codes of items.
In conclusion, each component of UNGER is essential to enhancing the effectiveness of generative recommendation.
@@
4.5. Scaling Law Study (RQ4)
Generative models have gained considerable attention due to their potential to exhibit scaling law properties. However, research exploring this phenomenon in the context of generative recommendation remains limited. Although recent study like HSTU (Zhai et al., 2024) has examined its scaling law characteristics, they are based solely on collaborative information derived from pure item IDs. To the best of our knowledge, no studies have investigated the scaling law of models that integrate both semantic and collaborative information. In this work, we are the first to explore this and compare our method with HSTU to analyze the scaling law differences between the two approaches.
As shown in Figure 6, while HSTU demonstrates scaling law on the MovieLens-20M dataset in (Zhai et al., 2024), it exhibits performance stagnation on the Amazon Beauty dataset once the model reaches eight layers. Beyond this point, the model’s performance begins to decline as the number of layers increases, a trend consistent with the observations in Table 1 of (Guo et al., 2024) on the Amazon Book dataset. In contrast, our method continues to show performance improvements on the Beauty dataset as the number of layers increases. Furthermore, our method consistently outperforms HSTU across all configurations, achieving a significant improvement of up to 21.42% in NDCG@20, demonstrating its superiority and effectiveness.
We attribute the superior performance of our method to its ability to simultaneously incorporate both collaborative and semantic knowledge with a unified code, effectively capturing their complementary nature. As a result, deeper models demonstrate enhanced capability in integrating these two types of knowledge, thus leading to consistent performance improvements.
@@ 4.6. Hyper-Parameter Analysis (RQ4)
Choice of different semantic encoders. To further investigate the influence of different pre-trained semantic encoders on recommendation performance, we conducted additional experiments on the Beauty dataset. Specifically, we employed various pre-trained semantic encoders to extract item semantic knowledge and evaluated their effectiveness using our proposed method. As presented in Figure 9, the results reveal the following insights: (1) Llama2-7b achieved the best performance, likely due to its substantially larger training corpus compared to Sentence-T5 and BERT. This extensive corpus enables Llama2-7b to produce richer semantic embeddings, which allow the unicodes to capture more comprehensive semantic knowledge, thereby enhancing the model’s ability to perform generative recommendation. (2) Our method is inherently plug-and-play, meaning it can directly incorporate the latest advancements from large language models without requiring any modifications to the underlying model architecture. This flexibility allows us to fully leverage the capability of state-of-the-art large language models, leading to improved recommendation performance.
Choice of different quantization methods. To evaluate the impact of different quantization methods on model performance, we conducted a comparative analysis on the Beauty dataset. We apply three different quantization methods: random assignment, RQVAE, and hierarchical K-means to discretize item representations. As illustrated in Figure 9, we observe that hierarchical K-means outperformed the other two methods. Compared to meaningless ID information, hierarchical K-means effectively captures the hierarchical relationship among items, where items sharing the same prefix codes are more similar than different prefix codes and thus more likely to be recommended to users. Additionally, in practice, we find that the RQVAE method suffers from persistent codebook conflicts during quantization, where multiple items are frequently assigned to a same code. To mitigate this issue, following a common practice (Rajput et al., 2023), we append a unique identifier at the end to ensure uniqueness. However, this approach partially disrupts the inherent hierarchical structure of the codebook, which could result in suboptimal performance in downstream recommendation tasks.
Number of Clusters. We analyzed the effect of the number of clusters $K$ in the hierarchical K-means algorithm on model performance on the Beauty dataset. Specifically, we vary $K$ across 128, 256, and 512. As shown in Figure 9, increasing $K$ from 128 to 256 enhances model performance, likely because a smaller $K$ limits the representational capacity of the codebook, making it hard to assign a distinct unicode to each item. Consequently, many dissimilar items are grouped under similar codes, reducing representational accuracy. When $K$ increases further from 256 to 512, the NDCG@20 metric improves, but the Recall@20 metric declines. This discrepancy may arise due to the larger search space, which increases the complexity of the decoding process. These findings highlight that for generative recommendation systems, selecting an appropriate number of clusters proportional to the total number of items is essential to balancing representation capacity and computational complexity.

### v3

@@ 4.2. Performance Comparison (RQ1)
4.2.1. Baselines
We compare UNGER with several representative related baselines, encompassing classical sequential modeling approaches as well as the latest emergent generative techniques.
@@ 4.2.1. Baselines
  • ColaRec (Wang et al., 2024b) integrates user-item interactions and content data within an end-to-end framework, leveraging pretrained collaborative identifiers, an item indexing task, and contrastive loss to align semantic and collaborative spaces for enhanced recommendation performance.
@@
  • EAGER (Wang et al., 2024c) integrates behavioral and semantic information through a two-stream architecture with shared encoding, separate decoding, and strategies for contrastive and semantic-guided learning to enhance collaborative information utilization.
4.2.2. Results
@@
Table 3 reports the overall performance of three datasets. The results for all baselines without the superscript are taken from the publicly accessible results (Zhou et al., 2020; Wang et al., 2024c). For missing statistics, we reimplement the baseline and report our experimental results. From the results, we have the following observations:
  • UNGER consistently achieves superior performance compared to base models across multiple datasets. Specifically, on the Beauty dataset, our proposed model UNGER outperforms the second-best baseline, EAGER, by a substantial margin, achieving improvements of 14.68% in Recall@20 and 7.85% in NDCG@20. These results underscore the effectiveness of our approach in real-world recommendation settings. We attribute this significant performance gain to UNGER’s novel ability to effectively integrate the complementary strengths of collaborative filtering signals and semantic content representations through a unified discrete code. By employing a compact representation space, UNGER not only enhances model efficiency but also reduces redundancy, achieving better performance with approximately half the number of total parameters compared to EAGER. This makes UNGER not only more accurate but also more lightweight and scalable.
  • UNGER surpasses prior generative recommendation models on most datasets. In head-to-head comparisons with previous generative baselines, UNGER exhibits a consistent advantage in performance metrics. This superiority stems from our model’s two-stage multi-task joint optimization framework, which is specifically designed to capture and unify collaborative and semantic knowledge within a shared code space. This framework enables UNGER to learn more coherent and expressive item representations by reinforcing the mutual information between collaborative signals and semantic embeddings. The unified code serves as a bridge that tightly couples these two modalities, thereby allowing the model to make more informed and nuanced recommendations. These improvements affirm the validity of our design and emphasize the importance of unified code representations in generative recommendation models.
  • Generative models generally outperform classical recommendation models in most cases. Our results reveals a clear trend: generative models tend to outperform traditional approaches such as sequential models based purely on item IDs. A key differentiator lies in how items are represented. Traditional methods often treat item IDs as atomic, non-informative symbols, lacking any structural or contextual information. In contrast, generative models, especially those based on Transformer decoders, leverage discrete hierarchical codes that encapsulate rich prior knowledge about item semantics and relationships. This structured representation aligns naturally with the token-by-token generation process of Transformers, enabling the model to better capture sequential dependencies and contextual relevance. As a result, generative models not only deliver better accuracy but also provide a more interpretable and principled framework for recommendation generation.
@@
Figure 6. T-SNE visualization of different embeddings distribution after applying K-means clustering (clusters=10).:
@@ 4.3. Semantic Domination Issue (RQ2)
To investigate whether our method addresses the semantic dominance issue, we conduct experiments on the Amazon Beauty dataset. We first extract the semantic embeddings of the items, then use only the item IDs from historical interaction data and apply the classical DIN model to obtain the collaborative embeddings. Following the most widely used vanilla approach of fusing knowledge from different modalities (Peng et al., 2022), concatenation, we concatenate the semantic embeddings with the collaborative embeddings. Notably, to mitigate the potential issues arising from mismatches in both dimensionality and value distribution between the two modalities, we first apply PCA to reduce the dimension of the semantic embeddings to match that of the collaborative embeddings. Then, we apply normalization to both embeddings before concatenating them to obtain the final concatenation embeddings. In addition, we obtain the integrated embeddings of items using our proposed method. We use T-SNE to visualize the distribution of these different embeddings. To more clearly illustrate the differences in their distributions, we apply K-means clustering algorithm to these embeddings, using different colors to represent data points from each cluster, as shown in Figure 6.
@@
As shown in Table 6, the concatenation method yields notably different similarity proportions between the semantic and collaborative modalities. Figure 6 further reveals that the resulting fused embeddings share a distributional pattern closely aligned with the original semantic embeddings. We refer to this phenomenon as the semantic dominance problem, which we define as a distributional imbalance where semantic modality disproportionately influences the final representation, thereby diminishing the contribution of the collaborative modality. While the two modalities are embedded in spaces of equal dimensionality and similar value ranges, their intrinsic data characteristics (e.g., semantic embeddings often being pretrained on large-scale textual data with smooth distributional properties) can lead to asymmetric influence when directly concatenated. Concatenation method tends to overly emphasize semantic knowledge while greatly ignoring collaborative knowledge, which poses a substantial obstacle to downstream recommendation effectiveness. In contrast, our proposed method effectively addresses this issue. It not only learns embeddings from both collaborative and semantic modalities in a more balanced manner at the distributional level but also achieves a more proportional similarity from each modality in quantitative evaluations. This prevents the semantic embeddings from dominating the final representation. As illustrated in Table 7, our method significantly outperforms both single-modality approaches and the concatenation-based fusion method, achieving the best results across all evaluation metrics. Notably, the concatenation method even performs worse than the semantic-only approach, indicating a negative impact on performance. This further highlights the severity of the semantic dominance problem and underscores the advantages of our proposed approach. We attribute these improvements to the successful integration of collaborative and semantic knowledge, which, unlike the concatenation method, better captures the complementary nature of the two different knowledge, enabling a more balanced and effective fusion of information from both modalities.
@@ 4.4. Ablation Study (RQ3)
To better understand the individual contributions of the components in our proposed model, UNGER, we conducted a comprehensive ablation study. Specifically, we incrementally removed two key auxiliary tasks, intra-modality knowledge distillation (IKD) task and cross-modality knowledge alignment (CKA) task to construct several ablated versions of the model. The corresponding performance of each variant is reported in Table 8.
  • Impact of Removing IKD and CKA. We observe that removing either the CKA or IKD task leads to a noticeable degradation in performance. When both tasks are excluded, the model reduces to a basic architecture where semantic and collaborative embeddings are merely concatenated. This baseline yields the poorest performance among all variants. These results strongly suggest that both IKD and CKA play indispensable roles in the success of UNGER.
  • Effectiveness of the CKA Task. When only the CKA task is retained, the model already exhibits significant performance gains over the baseline. This demonstrates that aligning knowledge across collaborative and semantic modalities allows the model to capture richer and more holistic item representations. The CKA task serves as a bridge between modalities, allowing the model to harness their complementary information effectively.
  • Synergistic Effect of CKA and IKD. The best performance is achieved when both the CKA and IKD tasks are jointly incorporated. In this setting, CKA establishes strong cross-modal connections while IKD reinforces intra-modal coherence, enabling the model to fully exploit the complementary strengths of collaborative and semantic knowledge. This synergy facilitates the learning of powerful, unified item codes that significantly enhance recommendation quality.
In conclusion, the results of the ablation study clearly demonstrate that each component of UNGER contributes meaningfully to its overall performance. The IKD and CKA tasks are not only effective in isolation but also highly complementary. Their integration is essential to achieving state-of-the-art results in generative recommendation, validating the effectiveness and superiority of our method.
Figure 7. Impact of the number of model layers on the Beauty dataset. The number of layers varies from 1 to 32.:
4.5. Scaling Law Study (RQ4)
Generative models have gained considerable attention due to their potential to exhibit scaling law properties. However, research exploring this phenomenon in the context of generative recommendation remains limited. Although a recent study, HSTU (Zhai et al., 2024), has investigated the scaling law characteristics of generative recommendation, it is restricted to collaborative signals derived purely from item IDs and only examines the impact of model depth. This narrow focus limits a more comprehensive understanding of scaling behavior. In contrast, to the best of our knowledge, no prior studies have explored scaling laws in models that jointly leverage both semantic and collaborative information. In this work, we make an initial and systematic attempt to explore the scaling behavior of generative recommendation models, examining how model performance evolves with respect to model depth, model width, and data volume.
4.5.1. Model Depth
To investigate the effect of model depth on generative recommendation performance, we vary the number of layers from 1 to 32 and evaluate the models on the Beauty dataset. As illustrated in Figure 7, the overall trend indicates consistent performance improvement as the model layers increases. Both Recall@20 and NDCG@20 show steady gains, particularly within the range of 1 to 8 layers, demonstrating the increasing capacity of deeper models to capture complex user-item interaction patterns.
We attribute the superior performance of our method to the unified code structure, which jointly encodes collaborative and semantic information into a shared representation space. This unified design enables the model to leverage both low-level interaction signals and high-level semantic attributes, thus offering richer and more diverse learning signals at every layer. In addition, the unified code structure mitigates redundancy between modalities and facilitates mutual regularization, allowing the model to disentangle meaningful patterns more effectively as it scales.
This synergy becomes increasingly beneficial with greater model depth. Deeper layers are able to refine and abstract over both types of information in a complementary manner, enhancing the model’s ability to generalize without overfitting to a single modality. Consequently, our full model consistently outperforms its ablated variants across all depth settings, and the performance gap widens as depth increases, highlighting the advantage of integrating both semantic and collaborative modalities in a unified framework.
Figure 8. Impact of data volume on the Beauty dataset.:
Figure 9. Impact of model width on the Beauty dataset. The number of width varies from 24 to 768.:
4.5.2. Data Size
We further examine how data volume influences the performance of generative recommendation by training models on subsets of the Beauty dataset, ranging from 20% to 100% of the full data. As shown in Figure 8, both Recall@20 and NDCG@20 consistently improve with increasing data size, suggesting that the model effectively benefits from larger training corpora.
This upward trend underscores the data-efficient nature of our unified code structure, which jointly leverages collaborative and semantic modalities. By learning from both interaction histories and item semantics, the model is able to extract more meaningful patterns as the data scale increases. In particular, Recall@20 experiences a significant gain when the data volume increases from 20% to 40%, suggesting that the model can rapidly leverage additional interaction signals to enhance its effectiveness. Beyond this point, model performance continues to improve steadily with more data, highlighting the scalability of our approach. The consistent upward trend across all data volumes demonstrates the importance of training data scale in enhancing the capacity of generative recommendation models.
4.5.3. Model Width
Besides, We also investigate how model width influences the performance of generative recommendation by varying the embedding dimension from 24 to 768. As presented in Figure 9, both Recall@20 and NDCG@20 exhibit an initial upward trend, followed by a performance decline as the width continues to increase.
The model achieves the best performance when the width is set to 96, indicating that a moderate level of width provides a good balance between model capacity and generalization. Increasing the width beyond this point does not lead to further improvement and even results in decreased performance, likely due to over-parameterization and reduced training stability under limited data conditions.
These results highlight the importance of selecting an appropriate width to fully utilize the benefits of our unified code structure. While wider models offer higher representational capacity, overly large widths may introduce optimization difficulties and lead to suboptimal generalization. Therefore, a carefully tuned width is essential for generative recommendation.
@@ 4.6. Hyper-Parameter Analysis (RQ4)
4.6.1. Choice of different semantic encoders
To investigate the influence of different pretrained semantic encoders on recommendation performance, we conduct a comprehensive comparative analysis on the Beauty dataset. Specifically, we explore three representative language models for extracting item-level textual representations: (1) BERT (Devlin et al., 2019), a bidirectional transformer trained with a masked language modeling objective, commonly used for general-purpose language understanding tasks; (2) Sentence-T5 (Ni et al., 2021), an encoder-decoder model optimized for semantic similarity and text matching, particularly effective on short-text tasks; and (3) Llama2-7b (Touvron et al., 2023), a large-scale decoder-only model trained on a diverse and extensive corpus, capable of capturing rich contextual and domain-specific semantics. These language models serve as plug-and-play components within our framework, allowing us to isolate and compare their impact on downstream recommendation quality under a unified training and evaluation protocol.
As presented in Figure 10, the results reveal the following insights: (1) Llama2-7b achieved the best performance, likely due to its substantially larger and more diverse training corpus, which covers a wide range of domains including news articles, encyclopedic content, social media posts, and source code. This extensive and heterogeneous data exposure enables Llama2-7b to learn richer contextual semantic relationships and finer-grained concept representations. In contrast, Sentence-T5, although specifically optimized for short-text semantic matching, is primarily trained on translation and summarization tasks. As a result, it struggles to fully capture the diverse, domain-specific, and stylistically nuanced nature of e-commerce product descriptions, which often involve specialized terminology and dynamic fashion-related expressions. BERT, while known for its stability on general-purpose language understanding tasks, employs a bidirectional masked language modeling objective that is not explicitly designed for semantic generation. Consequently, its embeddings are less effective in distinguishing polysemous words, capturing long-tail attributes, and modeling fine product-level distinctions. These limitations reduce the effectiveness of downstream residual quantization and semantic transfer in the generative recommendation stage, leading to suboptimal performance. (2) Our method is inherently plug-and-play, meaning that it allows the semantic encoder to be replaced without altering the residual quantization mechanism or the generative decoder. This design makes it easy to integrate the latest advancements in language modeling, such as DeepSeek-R1 (Guo et al., 2025), GPT-4o (Hurst et al., 2024), or domain-specific encoders trained on vertical data like fashion or electronics. Furthermore, our method can be extended to incorporate multimodal encoders that combine textual and visual inputs, such as CLIP-like models (Radford et al., 2021; Liu et al., 2023; Li et al., 2022), as long as the output dimensions are compatible. This modular and extensible architecture ensures that our framework can continually benefit from progress in pretrained representation learning, leading to consistent improvements in recommendation performance across domains.
4.6.2. Choice of different quantization methods.
To evaluate the impact of different quantization strategies on model performance, we conduct a comprehensive comparative analysis on the Beauty dataset. Specifically, we consider three distinct methods for discretizing continuous item representations: (1) Random Assignment, where items are randomly mapped to discrete codes without considering semantic or structural relationships; (2) RQVAE (Lee et al., 2022), which performs multi-level residual vector quantization through deep learning to capture high-dimensional semantics; and (3) Hierarchical K-means, a tree-structured clustering method that recursively partitions the embedding space to preserve hierarchical relations among items.
As illustrated in Figure 11, the hierarchical K-means approach significantly outperforms the other two methods across various evaluation metrics. This result highlights the importance of preserving structural information during the quantization process. In particular, hierarchical K-means encodes items such that those sharing common prefix paths in the hierarchical tree tend to be semantically or functionally similar. This structure enables the model to exploit hierarchical proximity during recommendation, improving the likelihood of retrieving items relevant to user preferences.
In contrast, the random assignment method fails to provide any meaningful structure to the discretized space, resulting in a noticeable drop in performance. The codes in this method do not reflect any semantic similarity, thereby limiting the model’s ability to generalize based on code proximity. While the RQVAE method is theoretically capable of modeling complex item distributions through its residual quantization mechanism, in practice we observe that it frequently suffers from codebook collisions. This refers to the phenomenon where multiple semantically distinct items are assigned to the same code due to limited codebook diversity or suboptimal training dynamics. Such collisions reduce the representational capacity of the quantized space and negatively impact recommendation performance. To address this issue, we adopt a mitigation strategy inspired by prior work (Rajput et al., 2023), wherein a unique identifier is appended to the learned code to enforce item-level uniqueness. Although this heuristic alleviates the collision problem, it compromises the learned hierarchical structure, as the appended identifier does not reflect semantic similarity. Consequently, this workaround introduces noise into the code space and may lead to suboptimal downstream performance.
Overall, these observations collectively underscore the advantages of hierarchical quantization methods in capturing and preserving the structural and semantic relationships among items, which are crucial for effective recommendation.
4.6.3. Number of Clusters.
We conduct an empirical investigation into how the number of clusters $K$ in the hierarchical K-means algorithm influences the overall performance of our method on the Beauty dataset. In particular, we experiment with three different cluster sizes: $K=128$, $K=256$, and $K=512$. The results are summarized in Figure 12.
Our analysis reveals that increasing $K$ from 128 to 256 leads to a consistent improvement in both Recall@20 and NDCG@20 metrics. This improvement is likely due to the enhanced representational capacity of the codebook as the number of clusters increases. When $K$ is relatively small (e.g., 128), the limited number of cluster centroids restricts the ability of the model to assign unique or highly discriminative codes to individual items. As a result, semantically or behaviorally dissimilar items may be forced to share similar codes, thereby reducing the granularity and accuracy of the learned item representations.
However, when we further increase $K$ from 256 to 512, we observe a more nuanced outcome. Specifically, the NDCG@20 metric continues to improve, indicating that the ranking quality of the top predicted items benefits from the more expressive codebook. This suggests that the model becomes better at placing the truly relevant items at higher ranks in the recommendation list. On the other hand, the Recall@20 metric slightly declines in this setting. We hypothesize that this is due to the expanded search space introduced by a larger $K$, which increases the decoding complexity and may lead to suboptimal retrieval results during inference. In other words, while the model has more representational freedom, the difficulty of accurately decoding from a much larger codebook may offset some of the gains.
Anyway, these findings collectively suggest that in generative recommendation systems, the number of clusters $K$ used in hierarchical K-means should be chosen carefully to strike a balance between representational richness and computational efficiency. A value of $K$ that is too small can constrain the model’s ability to encode item uniqueness, while an excessively large $K$ can introduce challenges in decoding and generalization. Selecting a cluster size proportional to the overall number of items in the dataset appears to be a practical heuristic for achieving this balance.
Figure 13. Sensitivity of hyper-parameter $\alpha$
Figure 14. Sensitivity of hyper-parameter $\beta$
4.6.4. Sensitivity of hyper-parameter $\alpha$
We investigate the sensitivity of the hyper-parameter $\alpha$ in the equation (5), which balances the next-item prediction and cross-modality knowledge alignment loss. As shown in Figure 13, we vary $\alpha$ from 0.2 to 5 and observe its effect on model performance.
Both Recall@20 and NDCG@20 improve significantly when $\alpha$ increases from 0.2 to 1.0, indicating that incorporating the alignment loss is beneficial for learning unified representations that integrate both collaborative and semantic information. Performance peaks when $\alpha$ is set to 1.0 or 2.0, suggesting that moderate emphasis on the alignment objective leads to optimal integration of modality-specific knowledge.
Importantly, the performance remains stable across a broad range of $\alpha$ values (from 0.5 to 5.0), demonstrating the robustness of the proposed method to this hyper-parameter. This insensitivity implies that our model is not overly reliant on fine-tuning $\alpha$, and can consistently benefit from the alignment signal as long as it is sufficiently incorporated. This robustness is particularly desirable in practical scenarios where extensive hyper-parameter search may be computationally prohibitive.
Figure 15. An example that illustrates the application of UNGER for the video game recommendation scenario.
4.6.5. Sensitivity of hyper-parameter $\beta$
We analyze the sensitivity of the hyper-parameter $\beta$ used in the equation (12), which balances the generative recommendation objective and the intra-modality knowledge distillation task. As illustrated in Figure 14, we vary $\beta$ from 0.2 to 5.0 and evaluate the model performance.
Both Recall@20 and NDCG@20 initially improve as $\beta$ increases from 0.2 to 1.0, demonstrating the benefit of incorporating the distillation loss for enhancing the quality of the learned representations. This confirms that the global semantic signal distilled via the learnable token $c_{\text{dis}}$ provides complementary supervision that helps compensate for the information loss caused by quantization in Stage I.
The model achieves its best performance when $\beta=1.0$, but remains stable across a relatively wide range (from 0.5 to 2.0), indicating that the method is robust to the choice of this hyper-parameter. Even when $\beta$ is increased to 5.0, performance does not deteriorate drastically, suggesting that the optimization process can effectively balance the two objectives. This robustness is particularly useful in practice, as it reduces the reliance on extensive hyper-parameter tuning during Stage II training.
4.7. Case Study
To demonstrate the interpretability and effectiveness of the proposed UNGER generative recommendation framework, we present a case study centered on personalized video game recommendation as shown in Figure 15. This study highlights how UNGER performs autoregressive decoding over discrete unified codes, which jointly encode semantic, behavioral, and contextual information to infer user intent and generate recommendations that align with multifaceted user preferences.
4.7.1. User Historical Interaction Profiling
We consider a representative user whose interaction history includes the following video games: Black Myth: Wukong, Cyberpunk 2077, Assassin’s Creed Shadows, and Call of Duty: Black Ops. These titles span multiple genres, narratives, and production styles. However, a pattern emerges when their key unified features are examined. The unified codes extracted from these items reflect consistent user affinities across several dimensions:
  1. (1)
    Gameplay Type and Genre: Action-oriented experiences that involve combat, stealth, and open-world exploration.
  2. (2)
    Audience Targeting: Mature content featuring violence, morally complex narratives, and sophisticated storytelling.
  3. (3)
    Platform Engagement: A clear preference for PlayStation-compatible titles, suggesting a stable user-platform context.
Unlike purely semantic models, UNGER does not rely on handcrafted or static attributes. Instead, each historical item is transformed into a sequence of discrete unified codes, learned through the training process to encapsulate relevant item-level and user-dependent information. These codes serve as the foundation for personalized sequence modeling.
4.7.2. Autoregressive Unified Code Decoding
Based on the unified codes derived from the user’s interaction history, UNGER proceeds with autoregressive decoding to infer the next most probable unified code sequence: $[66,100,168,99]$. We first conduct a clustering analysis between the code distribution and item attributes, through which we find that certain codes consistently correspond to specific semantic concepts. Each code thus represents a learned latent concept that captures diverse patterns, such as item semantics, user intent, or platform preference. The decoding trajectory can be interpreted as follows:
  1. (1)
    Code 66 (Domain: Video Games). Anchors the generation within the gaming domain, excluding items from other verticals such as music or movies.
  2. (2)
    Code 100 (Mature Content Affinity). Indicates the model’s inference of a mature content preference, derived from recurring patterns in prior game choices.
  3. (3)
    Code 168 (Action-Oriented Gameplay). Captures the user’s interest in fast-paced, action-heavy interactions involving combat, tactics, or physical exploration.
  4. (4)
    Code 99 (Platform: PlayStation). Implies a behavioral pattern involving consistent engagement with PlayStation-specific or compatible titles.
While each code does not correspond directly to a human-defined label, their learned representations align with high-level concepts, allowing for interpretable projections. The unified code space thereby serves as a latent reasoning structure that guides generation.
4.7.3. Recommendation Interpretation: Red Dead Redemption
The final recommendation produced by UNGER is Red Dead Redemption. We map the generated unified code sequence to the profile of this title and find a high degree of alignment:
  • Code 66 (Video Game Domain): Red Dead Redemption is firmly situated in the video game domain.
  • Code 100 (Mature Content): The game’s ESRB rating is Mature, consistent with the user’s demonstrated comfort with adult themes.
  • Code 168 (Action-Adventure Gameplay): The gameplay combines narrative, exploration, and action elements, mirroring previous interaction patterns.
  • Code 99 (Platform): Available on PlayStation, the game respects the user’s device and ecosystem usage pattern.
Despite the fact that unified codes are not manually engineered semantic tags, their learned representations align well with interpretable item features, thereby enabling natural explanations post-hoc.
4.7.4. Discussion and Insights
This case study highlights UNGER’s capability to bridge implicit user behavior and explicit recommendation reasoning via a discrete code generation process. Compared to traditional recommendation systems that rely on uninterpretable latent embeddings or static feature matching, UNGER leverages unified code modeling to incrementally build a high-fidelity representation of user intent. The sequential decoding procedure refines the search space across multiple dimensions, which include domain, content style, gameplay genre, and platform, thus mimicking a step-by-step decision-making process similar to how users articulate preferences.
Moreover, the discrete and interpretable nature of the unified code space provides a transparent layer between user history and item selection. The recommendation of Red Dead Redemption exemplifies how UNGER can reconstruct complex, multi-attribute user intent and map it to appropriate content, validating the framework’s design objectives: transparency, personalization, and controllable generation.